• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Jixie Zhang

    Log entry time 16:13:19 on July10,2010

    Entry number 331342

    keyword=Day shift summary

    Events Summary:
    08:00 Shift start with beam off. We are waiting for Jack to fix the right Q1
    08:30 Called MCC for escort access, will change to controlled access while Jack working in the hall
    08:50 Hall was now in controlled access
    12:00 Jack was out. Go back to beam permission.
    13:00 Spot++ test for run 2010
    13:34 Changed laser A and B mmixtrue to higher beam current
    13:58 Did test run with T1 prescales calibration. Run 2016 is for T1 calibration
    14:19 2018 Q1 tripped
    14:21 Took run 2019 for T3 calibration.
    14:30 We found that the the field of right Q1 was not reset to the right value. Before fixed it is 0.01414 and after is 0.01402. Now we recovered the field to 0.1414. Turn the regulation of right Q1 off.
    Runs from 2010 to 2020 may be affected.
    14:58 Run 2022 had zero rate. Tried to end it but failed. DAQ crashed! Reboot the QAQ. Both runs 2022 and 2023 are junk.
    15:15 MCC called and told that they can sent us 150 micro A beam. We will go for it.
    15:40 It seemed the beam current was not stable at 150 micro A, asked MCC to change to what ever high current they can go.
    15:55 Increased the field of right Q1 to 0.1416, with current 563.395 A.
    16:00 End the shift with data taking with run 2027.


    Runs Summary:
    Run T6Counts Comment
    2010 spot++ test
    2011 19.8K
    2012 21.5K
    2013 4.2K
    2014 21.7K
    2015 21.5K
    2016 T1 Calibration
    2017 junk
    2018 junk
    2019 T3 Calibration with T6
    2020 junk
    2021 21.4K
    2022 junk
    2023 junk
    2024 10.2K
    2025 20.7K
    2026 11.3K